Hi! I need some help with this problem and there is 3 parts to it.
#1. Is f(x)=3x-3 a function?
#2. Evaluate f(a+1) for f defined in problem #1.
#3. What are the domain and range for f defined in problem #1?
Thanks so much for your help!

Answer by funmath(2873)

#1. Is f(x)=3x-3 a function?
Yes. For every value of x, there is one and only one f(x).
:
#2. Evaluate f(a+1) for f defined in problem #1.
Substitute a+1 for x.
f(a+1)=3(a+1)-3
f(a+1)=3a+3-3
f(a+1)=3a
:
#3. What are the domain and range for f defined in problem #1?
There are no restrictions on x in a polynomial, so the domain is ALL REAL NUMBERS. in set builder notation: {x|x=R}
Because x can range from negative infinity to positive infinity, the range is also ALL REAL NUMBERS. in set builder notation: {y|y=R}
:
Happy Calculating!!!